# DiaBlo: Diagonal Blocks Are Sufficient For Finetuning

### v1

@@ 3.2 Comparison with LoRA
These gradients depend on the values of both $\mathbf{A}$ and $\mathbf{B}$, and can suffer from vanishing or unstable updates. As a result, LoRA’s performance is sensitive to initialization strategies and often requires carefully designed optimization techniques. In contrast, DiaBlo directly applies gradients to the full-rank diagonal blocks, making it inherently more stable and effective for fine-tuning.
Low-Rank vs. Full-Rank
Although the block diagonal adaptation $\mathbf{D}$ in DiaBlo is initialized to zero, it is not constrained to be low-rank during training. In contrast to LoRA, which limits updates to a low-rank subspace, DiaBlo operates in a restricted full-rank subspace formed by the diagonal blocks of the weight matrix. While it is theoretically difficult to determine which subspace—low-rank or diagonal-block—is universally more expressive or optimal for fine-tuning, our empirical results suggest that the diagonal-block subspace used by DiaBlo is sufficiently expressive for a wide range of downstream tasks. In practice, DiaBlo consistently delivers strong performance, often surpassing LoRA in accuracy while maintaining similar efficiency.
@@ 4 Experiments
In this section, we evaluate the performance of DiaBlo on a variety of benchmark datasets. Specifically, we assess its effectiveness on commonsense reasoning tasks, arithmetic reasoning tasks, code generation tasks, and safety alignment tasks. Our experiments are conducted using LLaMA2-7B touvron2023llama2 , LLaMA3-8B Dubey2024TheL3 , and Mistral-7B-v0.1 Jiang2023Mistral7 models. All fine-tuning and evaluation are performed on a single NVIDIA A100 GPU with 80GB of memory. The implementation details are provided in Appendix A.
@@ 4.1 Commonsense Reasoning
Commonsense reasoning 170k includes eight benchmark datasets: BoolQ clark2019boolq , PIQA bisk2020piqa , SocialIQA sap2019socialiqa , HellaSwag zellers2019hellaswag , WinoGrande sakaguchi2021winogrande , ARC-Easy, ARC-Challenge clark2018think , and OpenBookQA mihaylov2018can . Each task is framed as a multiple-choice problem. We fine-tune the LLaMA2-7B and LLaMA3-8B models on the combined training data from all eight tasks, and report accuracy on the test set of each task individually.
Implementation Details
Our fine-tuning and evaluation follow the standard settings in hu-etal-2023-llm ; DoRA . Specifically, we fine-tune the Q, K, V, Up, and Down modules using a maximum sequence length of 256. The model is loaded in float16 precision, the DiaBlo adaptation is in float32 precision, and FP16 mixed precision training is enabled.
@@ Results
The results on Commonsense reasoning tasks are shown in Table 1. We compare DiaBlo with existing methods, including DoRA DoRA , Pissa Meng2024PiSSAPS , MiLoRA Wang2024MiLoRAHM , and SMT He2024SparseMI . DiaBlo consistently outperforms existing PEFT baselines on commonsense reasoning tasks using both Llama2-7B and Llama3-8B models in FP16 precision. For Llama2-7B, DiaBlo achieves the highest average score of 83.5% while using only 0.52% trainable parameters, surpassing all other methods with similar number of parameters by at least 1.7%. The SMT can reach an average of 83.4% but requires 4.91% trainable parameters. For Llama3-8B, DiaBlo with $N=64$ and 1.04% trainable parameters achieves the best overall average score of 87.3%, surpassing LoRA and DoRA by 6.5% and 2.1%, respectively. The best SMT variant achieves 87.2%, which uses significantly more parameters. DiaBlo also demonstrates strong and consistent results across individual tasks. The findings demonstrate DiaBlo is an efficient and effective fine-tuning strategy that offers competitive accuracy while using few trainable parameters.
@@
SMT He2024SparseMI proposes to select most important sub-matrices in pretrained weights based on magnitudes of gradients and then only fine-tune these sub-matrices. In contrast, DiaBlo directly selects diagonal blocks from the pretrained weights, offering a simpler and more implementation-friendly approach. The experiments further demonstrate that fine-tuning diagonal blocks provide even better results. As shown in Table 1, DiaBlo achieves an average score of 83.5% on Llama2-7B, significantly outperforming SMT’s 81.8% while using a less number of trainable parameters. Even when SMT increases its trainable parameter to 4.91% to achieve its best result, DiaBlo, with only 0.52% trainable parameters, still attains better performance. A similar trend holds for Llama3-8B, where DiaBlo outperforms SMT even when SMT uses considerably more trainable parameters. These results demonstrate that fine-tuning only the diagonal blocks of pretrained weights is not only highly efficient but also sufficient to achieve strong performance.
Precision
The experiments in existing literature, e.g., DoRA DoRA , Pissa Meng2024PiSSAPS , and MiLoRA Wang2024MiLoRAHM , all follow the fine-tuning setting in hu-etal-2023-llm , which uses FP16 for model weights, FP32 for low-rank adaptation, and FP16 mixed-precision training. Our results reported in Table 1 follow the same fine-tuning setting. However, we notice that LoRA finetuning in precision BF16 may achieve better results. For comparison, we also report results trained in BF16 precision following the setting in LoRI Zhang2025LoRIRC . The results are reported in Table 2. On LLaMA3-8B, DiaBlo achieves the highest average score of 88.3% with $N=64$ blocks and 1.51% of trainable parameters, surpassing LoRA, DoRA, and LoRI by at least 1.0%. Even when reducing the number of trainable parameters to 0.76% with $N=128$ blocks, DiaBlo maintains strong performance, achieving 88.0%, again outperforming the baselines. On Mistral-7B, DiaBlo achieves an average accuracy of 88.0% with $N=64$ blocks, exceeding LoRA, DoRA, and LoRI. At $N=128$ blocks with less trainable parameters, it continues to perform competitively, reaching 87.9%. These results confirm that DiaBlo is highly effective under BF16 precision, offering both superior accuracy and parameter efficiency compared to other state-of-the-art PEFT methods.
Table 1: Commonsense reasoning results of fine-tuning Llama2-7B and Llama3-8B in FP16. Results marked with , , are taken from Dora DoRA , MiLoRA Wang2024MiLoRAHM , and SMT He2024SparseMI , respectively.
@@
PEFT $\mathbf{r/N}$Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
LoRA $r=$32 0.83% 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6
DoRA $r=$32 0.83% 71.8 83.7 76.0 89.1 82.6 83.7 68.2 82.4 79.7
DoRA $r=$16 0.42% 72.0 83.1 79.9 89.1 83.0 84.5 71.0 81.2 80.5
Pissa $r=$32 0.83% 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA $r=$32 0.83% 67.6 83.8 80.1 88.2 82.0 82.8 68.8 80.6 79.2
@@
PEFT $\mathbf{r/N}$Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
LoRA $r=$32 0.78% 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
DoRA $r=$32 0.78% 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8 85.2
DoRA $r=$16 0.39% 74.5 88.8 80.3 95.5 84.7 90.1 79.1 87.2 85.0
Pissa $r=$32 0.78% 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA $r=$32 0.78% 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
@@
DiaBlo $N=$64 1.04% 75.8 90.8 80.7 95.6 89.9 93.4 83 89.2 87.3
DiaBlo $N=$128 0.52% 76.1 90.3 81.5 95.7 89.7 93.4 83.4 87.2 87.2
Table 2: Commonsense reasoning results of fine-tuning Llama3-8B and Mistral-7B in BF16 following LoRI. Results of baseline methods are taken from LoRI Zhang2025LoRIRC .
Mistral-7B
PEFT $\mathbf{r/N}$Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
LoRA $r=$32 1.25% 75.2 90.1 82.9 95.1 88.1 92.0 82.9 88.7 86.9
DoRA $r=$32 1.25% 75.8 90.4 82.9 96.3 87.9 92.6 83.3 90.6 87.5
LoRI $r=32$ 0.63% 75.9 90.6 83.0 95.9 87.4 91.9 83.6 88.4 87.1
DiaBlo $N=$64 1.68% 76.6 90.9 83.7 96.9 88.6 93.8 83.2 90.5 88.0
DiaBlo $N=$128 0.84% 76.7 91.3 83.9 96.9 88.2 93.1 83.4 89.7 87.9
Llama3-8B
PEFT $\mathbf{r/N}$Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
LoRA $r=$32 1.12% 76.3 89.8 82.7 95.8 88.7 91.7 83.4 88.4 87.1
DoRA $r=$32 1.12% 75.9 89.8 82.7 95.3 88.2 93.2 83.5 87.9 87.1
LoRI $r=32$ 0.56% 76.4 89.0 82.7 95.9 87.9 93.6 84.2 88.5 87.3
DiaBlo $N=$64 1.51% 76.6 90.2 83.1 96.9 88.8 93.9 85.1 91.5 88.3
DiaBlo $N=$128 0.76% 76.2 90.5 83.7 97.0 89.0 94.0 84.7 88.9 88.0
@@ 4.2 Arithmetic Reasoning
We evaluate DiaBlo on arithmetic reasoning using the MetaMathQA Yu2023MetaMathBY dataset, which consists of 395K samples augmented from the training sets of GSM8K cobbe2021training and MATH Hendrycks2021MeasuringMP . We fine-tune the LLaMA2-7B model on this dataset and evaluate performance on the test sets of GSM8K and MATH. The training follows the setting in MiLoRA and we report results from the final checkpoint and use the Exact Match (EM) ratio to measure accuracy.
@@ Results
On arithmetic reasoning tasks, DiaBlo demonstrates strong performance, achieving results that match or exceed full fine-tuning while maintaining parameter efficiency. With $N=32$ blocks and only 2.09% trainable parameters, DiaBlo attains an average accuracy of 43.4%, slightly outperforming full fine-tuning with 43.2% and significantly surpassing LoRA, PiSSA, and MiLoRA by 4.7%, 6.4%, and 2.7%, respectively. Even with $N=64$ blocks and 1.04% trainable parameters, DiaBlo with fewer trainable parameters maintains high performance with an average accuracy of 42.1%, demonstrating that it remains effective under restrictive parameter budgets. Notably, DiaBlo achieves the highest score of 20.4% on the MATH dataset among all methods, including full fine-tuning. These results highlight DiaBlo’s ability to generalize to complex reasoning tasks while offering superior parameter efficiency.
Table 3: Performance of fine-tuning Llama2-7B on arithmetic reasoning tasks. Results of baseline methods are taken from MiLoRA Wang2024MiLoRAHM .
Method $\mathbf{r}/\mathbf{N}$Params GSM8K MATH AVG
@@ 4.3 Code Generation and Safety Alignment
We further evaluate the performance of DiaBlo on code generation and safety alignment tasks, following the training setting in LoRI Zhang2025LoRIRC . For code generation, we fine-tune on the CodeAlpaca dataset codealpaca and evaluate using the HumanEval benchmark Chen2021EvaluatingLL with standard metrics: Pass@1, Pass@5, and Pass@10. For safety alignment, we fine-tune on the SaferPaca dataset Bianchi2023SafetyTunedLL , an extension of Alpaca-Cleaned alpaca , and assess performance based on the refusal rate to harmful prompts from the HEx-PHI dataset Qi2023FinetuningAL . We conduct experiments on both the LLaMA3-8B and Mistral-7B models.

### v2

@@ 3.2 Theoretical Analysis
These gradients depend on the values of both $\mathbf{A}$ and $\mathbf{B}$, and can suffer from vanishing or unstable updates. As a result, LoRA’s performance is sensitive to initialization strategies and often requires carefully designed optimization techniques. In contrast, DiaBlo directly applies gradients to the full-rank diagonal blocks, making it inherently more stable and effective for fine-tuning. The stability of DiaBlo is demonstrated by gradient norm comparisons with LoRA in Appendix A.3.4, where DiaBlo consistently exhibits lower variance, except for the vanishing $\mathbf{A}$-matrix gradients of LoRA in the early stage of fine-tuning.
DiaBlo Converges to A Solution of Full FT
In the linear least squares problem (LSQ) setting, DiaBlo is strictly more expressive than LoRA when the input matrix $\mathbf{X}$ has low rank. Consider the fine-tuning objective
$\min_{\mathbf{W}\in\mathbb{R}^{m_{1}\times m_{2}}}\frac{1}{2}\|\mathbf{Y}-\mathbf{X}\mathbf{W}_{0}-\mathbf{X}\mathbf{W}\|_{F}^{2},$ (3)
where $\mathbf{X}\in\mathbb{R}^{b\times m_{1}}$, $\mathbf{Y}\in\mathbb{R}^{b\times m_{2}}$, and $\mathbf{W}_{0}\in\mathbb{R}^{m_{1}\times m_{2}}$. For DiaBlo, we instead optimize over block-diagonal updates:
$\min_{\mathbf{D}=\mathrm{diag}(\mathbf{D}_{1},\ldots,\mathbf{D}_{N})}\frac{1}{2}\|\mathbf{Y}-\mathbf{X}\mathbf{W}_{0}-\mathbf{X}\mathbf{D}\|_{F}^{2},$ (4)
where $\mathbf{D}_{i}\in\mathbb{R}^{d_{1}\times d_{2}}$, $d_{1}=\frac{m_{1}}{N}$, and $d_{2}=\frac{m_{2}}{N}$. The following theorem shows that, under mild assumptions, any minimizer of the DiaBlo-LSQ (4) is also a minimizer of the full-LSQ (3).
Theorem 1.
Suppose that $\mathbf{X}$ is a generic rank-$r$ matrix. If the number of diagonal blocks $N\leq\frac{m_{1}}{r}$ is a common factor of $m_{1},m_{2}$, then any solution to the DiaBlo-LSQ (4) also solves the full-LSQ (3).
The proof of Theorem 1 is provided in Appendix A.1.1. According to the theorem, DiaBlo matches full FT performance with as few as $Nd_{1}d_{2}=\frac{m_{1}m_{2}}{N}\geq m_{2}r$ trainable parameters. In contrast, LoRA must use a rank of at least $r$ to solve the full-LSQ, necessitating at least $(m_{1}+m_{2})r$ parameters. Consequently, DiaBlo remains strictly more expressive than LoRA under the same parameter constraints.
Remarkably, DiaBlo’s effectiveness is not restricted to the linear setting. In general neural network fine-tuning, updating only diagonal blocks is often sufficient to match full FT performance. Our theoretical result substantiates this: DiaBlo converges to a stationary point of the full FT objective whenever the activation matrix $\mathbf{X}$ and output gradient $\mathbf{g}_{\mathbf{Y}}$ exhibit a low-rank—a property empirically observed in recent literature (Chee et al., 2023; Zhang et al., 2024; Zhao et al., 2024). Formally:
Theorem 2.
Suppose the activation matrix $\mathbf{X}\in\mathbb{R}^{b\times m_{1}}$ and the linear output gradient $\mathbf{g}_{\mathbf{Y}}\in\mathbb{R}^{b\times m_{2}}$ are generic with ranks $r_{1}$ and $r_{2}$, respectively. If $N$ is a common factor of $m_{1},m_{2}$ and satisfies $N\geq\lceil\frac{r_{1}N}{m_{1}}\rceil\lceil\frac{r_{2}N}{m_{2}}\rceil$, then any adaptation $\mathbf{D}=\mathrm{diag}(\mathbf{D}_{1},\ldots,\mathbf{D}_{N})\in\mathbb{R}^{m_{1}\times m_{2}}$ produced by DiaBlo that satisfies the stationarity conditions $\mathbf{g}_{\mathbf{D}_{i}}=\mathbf{0}_{d_{1}\times d_{2}}$ for all $1\leq i\leq N$ yields adapted weights $\mathbf{W}=\mathbf{W}_{0}+\mathbf{D}$ that is also a stationary point of the full FT objective, i.e., $\mathbf{g}_{\mathbf{W}}=\mathbf{0}_{m_{1}\times m_{2}}$.
The proof is provided in Appendix A.1.2. Ignoring the ceiling operators in the assumption $N\geq\lceil\frac{r_{1}N}{m_{1}}\rceil\lceil\frac{r_{2}N}{m_{2}}\rceil$ gives an approximate condition $N\leq\frac{m_{1}m_{2}}{r_{1}r_{2}}$. Consequently, lower ranks for $\mathbf{X}$ and $\mathbf{g}_{\mathbf{Y}}$ permit a larger numbers of blocks $N$ in Diablo. This reduces trainable parameters while maintaining the guarantee that the finetuned model is a stationary point of the full FT objective. An empirical validation of both the low-rank assumptions and the convergence of the full-weight gradient is presented in Appendix A.2.
@@ 4 Experiments
In this section, we evaluate the performance of DiaBlo on a variety of benchmark datasets. Specifically, we assess its effectiveness on commonsense reasoning tasks, arithmetic reasoning tasks, code generation tasks, and safety alignment tasks. Our experiments are conducted using LLaMA2-7B (Touvron et al., 2023b), LLaMA3-8B (Dubey et al., 2024), and Mistral-7B-v0.1 (Jiang et al., 2023) models. All fine-tuning and evaluation are performed on a single NVIDIA A100 GPU with 80GB of memory. Additional experiments are shown in Appendix A.3 and the implementation details are provided in Appendix A.4.
@@ 4.1 Commonsense Reasoning
Commonsense reasoning 170k includes eight benchmark datasets: BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SocialIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-Easy, ARC-Challenge (Clark et al., 2018), and OpenBookQA (Mihaylov et al., 2018). Each task is framed as a multiple-choice problem. We fine-tune the Llama2-7B, Llama3-8B, and Llama-13B models on the combined training data from all eight tasks, and report accuracy on the test set of each task individually.
@@ Results
The results on Commonsense reasoning tasks are shown in Table 1. We compare DiaBlo with existing methods, including DoRA (Liu et al., 2024), Pissa (Meng et al., 2024), MiLoRA (Wang et al., 2024a), and SMT (He et al., 2024). DiaBlo consistently outperforms existing PEFT baselines on commonsense reasoning tasks using Llama2-7B, Llama3-8B, and Llama-13B models in FP16 precision. For Llama2-7B, DiaBlo achieves the highest average score of 83.5% while using only 0.52% trainable parameters. For Llama3-8B, DiaBlo with $N=64$ achieves the best overall average score of 87.3%. For the larger Llama-13B model, DiaBlo again demonstrates superior performance with a score of 84.9%, significantly outperforming strong baselines. The findings demonstrate DiaBlo is an efficient, scalable, and effective fine-tuning strategy.
@@
SMT (He et al., 2024) proposes to select most important sub-matrices in pretrained weights based on magnitudes of gradients and then only fine-tune these sub-matrices. In contrast, DiaBlo directly selects diagonal blocks from the pretrained weights, offering a simpler and more implementation-friendly approach. The experiments further demonstrate that fine-tuning diagonal blocks provide even better results. As shown in Table 1, DiaBlo achieves an average score of 83.5% on Llama2-7B, significantly outperforming SMT’s 81.8% while using a less number of trainable parameters. Even when SMT increases its trainable parameter to 4.91% to achieve its best result, DiaBlo, with only 0.52% trainable parameters, still attains better performance. A similar trend holds for Llama3-8B, where DiaBlo outperforms SMT even when SMT uses considerably more trainable parameters. These results demonstrate that fine-tuning only the diagonal blocks of pretrained weights is not only highly efficient but also sufficient to achieve strong performance.
Table 1: Commonsense reasoning results of fine-tuning Llama2-7B, Llama3-8B, and Llama-13B in FP16. Results marked with , , are taken from Dora (Liu et al., 2024), MiLoRA (Wang et al., 2024a), and SMT (He et al., 2024), respectively.
@@
PEFT $\mathbf{r/N}$ #Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
Full FT N/A 100% 73.3 85.7 81 90.2 86.9 88.6 77.4 85.2 83.5
Zero-shot N/A 0% 61.9 47.4 2.2 12.3 0.0 12.5 10.0 23.4 21.2
LoRA r=r=32 0.83% 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6
DoRA r=r=32 0.83% 71.8 83.7 76.0 89.1 82.6 83.7 68.2 82.4 79.7
DoRA r=r=16 0.42% 72.0 83.1 79.9 89.1 83.0 84.5 71.0 81.2 80.5
Pissa r=r=32 0.83% 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA r=r=32 0.83% 67.6 83.8 80.1 88.2 82.0 82.8 68.8 80.6 79.2
@@
PEFT $\mathbf{r/N}$ #Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
Full FT N/A 100% 76.4 89.7 82.5 95.5 89.6 92.9 84.3 89.2 87.5
Zero-shot N/A 0% 56.3 66.3 32.2 25.2 26.6 24.1 21.5 24.8 34.6
LoRA r=r=32 0.78% 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
DoRA r=r=32 0.78% 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8 85.2
DoRA r=r=16 0.39% 74.5 88.8 80.3 95.5 84.7 90.1 79.1 87.2 85.0
Pissa r=r=32 0.78% 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA r=r=32 0.78% 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
@@
DiaBlo N=N=64 1.04% 75.8 90.8 80.7 95.6 89.9 93.4 83.0 89.2 87.3
DiaBlo N=N=128 0.52% 76.1 90.3 81.5 95.7 89.7 93.4 83.4 87.2 87.2
Llama-13B
PEFT $\mathbf{r/N}$ #Params BoolQ PiQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
Full FT N/A 100% 74.8 86.2 80.5 90.5 87.1 88.6 75.4 88.0 83.9
Zero-shot N/A 0% 59.5 47.4 16.0 22.9 0.3 15.8 12.0 24.0 24.7
LoRA $r=32$ 0.67% 72.1 83.5 80.5 90.5 83.7 82.8 68.3 82.4 80.5
DoRA $r=32$ 0.68% 72.5 85.3 79.9 90.1 82.9 82.7 69.7 83.6 80.8
DiaBlo $N=64$ 1.06% 74.7 86.9 81.2 91.5 87.4 89.8 78.2 90.2 84.9
DiaBlo $N=128$ 0.53% 75.7 84.8 82.0 90.6 86.8 90.1 77.1 89.2 84.5
@@ 4.2 Arithmetic Reasoning
We evaluate DiaBlo on arithmetic reasoning using the MetaMathQA (Yu et al., 2023) dataset, which consists of 395K samples augmented from the training sets of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). We fine-tune the LLaMA2-7B model on this dataset and evaluate performance on the test sets of GSM8K and MATH. The training follows the setting in MiLoRA and we report results from the final checkpoint and use the Exact Match (EM) ratio to measure accuracy.
@@ Results
As shown in Table 2, on arithmetic reasoning tasks, DiaBlo demonstrates strong performance, achieving results that match or exceed full fine-tuning while maintaining parameter efficiency. With $N=32$ blocks and only 2.09% trainable parameters, DiaBlo attains an average accuracy of 43.4%, slightly outperforming full fine-tuning with 43.2% and significantly surpassing LoRA, PiSSA, and MiLoRA by 4.7%, 6.4%, and 2.7%, respectively. Even with $N=64$ blocks and 1.04% trainable parameters, DiaBlo with fewer trainable parameters maintains high performance with an average accuracy of 42.1%, demonstrating that it remains effective under restrictive parameter budgets. Notably, DiaBlo achieves the highest score of 20.4% on the MATH dataset among all methods, including full fine-tuning.
Table 2: Performance of fine-tuning Llama2-7B on arithmetic reasoning tasks. Results of baseline methods are taken from MiLoRA (Wang et al., 2024a).
Method $\mathbf{r}/\mathbf{N}$ #Params GSM8K MATH AVG
@@ 4.3 Code Generation and Safety Alignment
We further evaluate the performance of DiaBlo on code generation and safety alignment tasks, following the training setting in LoRI (Zhang et al., 2025a). For code generation, we fine-tune on the CodeAlpaca dataset (Chaudhary, 2023) and evaluate using the HumanEval benchmark (Chen et al., 2021) with standard metrics: Pass@1, Pass@5, and Pass@10. For safety alignment, we fine-tune on the SaferPaca dataset (Bianchi et al., 2023), an extension of Alpaca-Cleaned (Taori et al., 2023), and assess performance based on the refusal rate to harmful prompts from the HEx-PHI dataset (Qi et al., 2023). We conduct experiments on both the LLaMA3-8B and Mistral-7B models.